# GraphCogent: Overcoming LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

### v1

@@ Appendix I LLM Prompts
4:Parse raw graph data via subgraph sampling
5:Transform text representations to adjacency lists using heuristic prompts
6:Verify transformation via Graph Verifier (quantity consistency + format compliance)
[…]
8:Buffer Module:
9:Construct complete graph from adjacency lists
10:Transform to multiple representations: NetworkX/PyG/NumPy
11:Build data indices (dimensionality, schema, metadata)
12:Store preprocessed graph data $\tilde{G}$
13:Execution Module:
14:if $T$ is in common toolset coverage then
15:  $\triangleright$ See Fig. 16 for in-toolset reasoning example
16:  Select appropriate tool from common toolset
17:  Retrieve required data format from Buffer Module
18:  Execute tool calling on $\tilde{G}$
19:  Return result $R$
20:else
21:  $\triangleright$ See Fig. 17 for out-toolset reasoning example
@@
23:  $\triangleright$ See Fig. 18 for model generation example
24:  Combine $M$ with $\tilde{G}$ from Buffer Module
25:  Execute generated model
26:  Return result $R$
27:end if
Figure 15: Example of Sensory Agent transformation process.
Figure 16: Example of Reasoning Agent’s in-toolset processing.
Figure 17: Example of Reasoning Agent’s out-toolset processing.
Figure 18: Example of Model Agent’s model generation.

### v2

@@ Appendix E GraphCogent Workflow
4:if $G$ is file-based (edgelist/txt/json) then
5:  Bypass transformation; forward $G$ directly to Buffer
6:else
7:  Apply sliding-window sampling (GEC-selected; default 50 edges with 5% overlap)
8:  Transform text to adjacency lists via heuristic prompts
9:  Verify via Graph Verifier (quantity consistency + format compliance)
10:end if
11:Buffer Module:
12:Construct complete graph from adjacency lists
13:Transform to multiple representations: NetworkX/PyG/NumPy
14:Build data indices (dimensionality, schema, metadata)
15:Store preprocessed graph data $\tilde{G}$
16:Execution Module:
17:if $T$ is in common toolset coverage then
18:  Select appropriate tool from common toolset
19:  Retrieve required data format from Buffer Module
20:  Execute tool calling on $\tilde{G}$; return $R$
21:else
@@
23:  Combine $M$ with $\tilde{G}$ from Buffer Module
24:  Execute generated model; return $R$
25:end if